# Partial Resilient Leader-Follower Consensus in Time-Varying Graphs

### v1

@@ I Introduction
Consensus is a process where multiple agents agree to a common value. However, standard consensus protocols are vulnerable to adversarial agents that transmit false information to the network, potentially causing significant performance degradation or even failure. This has motivated extensive research on resilient consensus, which ensures consensus in the presence of adversaries [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12].
Many works rely on Mean Subsequence Reduced (MSR)-type algorithms [13, 1, 3, 8]. In the algorithms, each non-adversarial (normal) agent discards some extreme values from its neighbors before updating its state, thereby ensuring that the updated value is not influenced by the adversaries. For many MSR-type algorithms to succeed, the communication graph is assumed to satisfy topological conditions called r-robustness or $(r,s)$-robustness, which are sufficient (and necessary) conditions for normal agents to achieve consensus within the convex hull of their initial values [1, 2].
A closely related problem is that of resilient leader-follower consensus, where a subset of agents (leaders) propagate a reference signal that the rest of the network (followers) aims to track, despite the presence of adversaries [14, 4, 15]. Resilient broadcasting problems have also been studied [16, 17, 2], where algorithms such as the Certified Propagation Algorithm (CPA) were proposed to enable trustworthy leader to reliably broadcast information despite the presence of faulty nodes. In [18], the authors addressed the problem of resilient distributed estimation by leveraging reliable agents that are directly connected to others. Additionally, [19] established conditions under which information can be resiliently transmitted from a set of source nodes to other nodes without direct access to the information.
@@
Despite these efforts, the system behavior under MSR algorithms in non-ideal settings — where these robustness conditions of the entire network fail — remains poorly understood. Some progress has been made in this direction for leaderless settings. For example,[26, 27] study community or cluster consensus, where multiple sets of agents converge to multiple distinct values when the robustness of the entire network is not sufficient. However, they also assume certain robustness within communities (clusters) and do not study cases where these conditions fail. In another line of work,[28] investigates the non-resiliency of a graph by examining the number of possible non-convergent nodes (agents that fail to reach consensus with any other nodes) when the network lacks sufficient robustness, but these results are limited to specific classes of graphs.
@@ II Preliminaries
We denote a simple time-varying digraph as $\mathcal{G}[t]=(\mathcal{V},\mathcal{E}[t])$ where $\mathcal{V}$ and $\mathcal{E}[t]$ are the finite vertex set and the time-varying directed edge set, respectively. We also denote $\mathbb{G}=(\mathcal{G}[t])_{t\in\mathbb{Z}_{\geq 0}}$ as a sequence of time-varying graphs. Within the set $\mathcal{V}$, we have leaders $\mathcal{L}=\{1,\dots,l\}\subset\mathcal{V}$ that propagate the reference state according to the same function $f_{r}:\mathbb{Z}_{\geq 0}\to\mathbb{R}$ to followers $\mathcal{F}=\{l+1,\dots,n\}=\mathcal{V}\setminus\mathcal{L}$. We denote $|\mathcal{L}|=l$ and $|\mathcal{F}|=f$. A directed edge $(i,j)\in\mathcal{E}[t]$ indicates that agent $j$ is able to receive information from agent $i$ at time $t$. This implies agent $i$ is an in-neighbor of agent $j$, and agent $j$ is an out-neighbor of agent $i$. The in-neighbor and out-neighbor sets of agent $i$ are denoted as $\mathcal{N}_{i}[t]=\{j\in\mathcal{V}\;|\;(j,i)\in\mathcal{E}[t]\}$ and $\mathcal{N}_{i}^{o}[t]=\{j\in\mathcal{V}\;|\;(i,j)\in\mathcal{E}[t]\}$, respectively. The extended in-neighbor set is given as $\mathcal{B}_{i}[t]=\mathcal{N}_{i}[t]\cup\{i\}$. We denote the cardinality of a set $\mathcal{S}$ as $|\mathcal{S}|$. We denote the set of non-negative and positive integers as $\mathbb{Z}_{\geq 0}$ and $\mathbb{Z}_{>0}$.
@@ II-A Bootstrap Percolation
At a fixed time step $t\in\mathbb{Z}_{\geq 0}$, bootstrap percolation (BP) models the spread of activation of nodes of $\mathcal{G}[t]$ from a set of initial active nodes $\mathcal{L}\subset\mathcal{V}$, given user-defined threshold $r\in\mathbb{Z}_{>0}$ [29]. In BP, each node is either active or inactive. At each iteration of BP, each agent $i\in\mathcal{V}$ becomes active if it has at least $r$ active neighbors and remains active until the process terminates, which happens after at most $f$ iterations.
@@
$q_{i}[t,k]=\begin{cases}1&\text{if $q_{i}[t,k-1]=1$},\\ 1&\text{if $\sum_{j\in\mathcal{N}_{i}[t]}q_{j}^{i}[t,k-1]\geq r$},\\ 0&\text{if $\sum_{j\in\mathcal{N}_{i}[t]}q_{j}^{i}[t,k-1]<r$},\end{cases}$ (1)
@@ II-B Resilient Consensus
We review resilient leader-follower consensus when agents cannot distinguish between leaders and followers. Let $x_{i}[t]\in\mathbb{R}$ denote the consensus state of agent $i$ at time $t$, and $x_{i}^{j}[t]$ the copy received by agent $j$. Each agent $i$ shares $x_{i}[t]$ with its out-neighbors and updates
@@
An agent $a\in\mathcal{V}$ is an adversary if it transmits consensus state $x_{a}^{u}[t]\in\mathbb{R}$ to all out-neighbors $u\in\mathcal{N}_{a}[t]$ for all $t\geq\mathbb{Z}_{\geq 0}$ but does not follow the update protocols (1) and/or (2), or if it does not send the same values to all out-neighbors at some time $t$.
The adversaries in this paper are a slight modification of the well-known Byzantine agents [1, 4]: each $a\in\mathcal{A}$ may manipulate both its consensus state $x_{a}^{u}[t]$ and activation state $q_{a}^{u}[t,k]$. Both followers and leaders can be adversaries. Non-adversarial agents are called normal agents. We denote $\mathcal{A}$ as the set of adversaries and $\mathcal{M}=\mathcal{V}\setminus\mathcal{A}$ as the set of normal agents, with $\mathcal{F}_{\mathcal{M}}$$\mathcal{F}\cap\mathcal{M}\neq\emptyset$ and $\mathcal{L}_{\mathcal{M}}$$\mathcal{L}\cap\mathcal{M}\neq\emptyset$.
@@
For all $a\in\mathcal{A}$, $u\in\mathcal{N}_{a}^{o}[t]$, $k\in\{0,\dots,f\}$, and $t\in\mathbb{Z}_{\geq 0}$, we have $q_{a}^{u}[t,k]\in\{0,1\}$. Furthermore, for any $a\in\mathcal{A}$, if $q_{a}^{u}[t,k^{\prime}]=1$ for some $u\in\mathcal{N}_{a}^{o}[t]$ at iteration $k^{\prime}<f$, then $q_{a}^{u}[t,k^{\prime\prime}]=1$ for all $k^{\prime\prime}\geq k^{\prime}$.
Hence, an adversary cannot (i) transmit non-binary values or (ii) switch $q_{a}^{u}[t,k]$ from $1$ to 0 during BP process. These are direct violations of (1) that would lead to detections, which adversaries would try to avoid. In addition, we consider the following scope of threat:
@@
Lemma 1 ([25, Corollary 1 (revised)]).
@@ III Problem Statement
Although adversaries may manipulate both their consensus and activation states, our focus is on the convergence of the consensus states. That is, we do not require the activation states received by a normal agent to be correct. Nevertheless, as we demonstrate in later sections, how adversaries share their activation states influence which followers ultimately achieve convergence (see Section IV-B).
@@
Let $\mathbb{G}=(\mathcal{G}[t])_{t\in\mathbb{Z}_{\geq 0}}$ be an arbitrary sequence of time-varying digraphs with an $F$-local $\mathcal{A}$. We aim to characterize sufficient conditions under which followers in $\mathcal{F}_{\mathcal{C}}\subseteq\mathcal{F}_{\mathcal{M}}$ achieve partial resilient leader–follower consensus, even when sufficient robustness conditions do not hold.
@@ IV Partial Resilient Leader-Follower Consensus
Consider a network $\mathcal{G}[t]=(\mathcal{V},\mathcal{E}[t])$ that fails to be strongly $r$-robust at time $t$. Since robustness conditions are defined over the entire network topology, that implies there exists a nonempty subset $\mathcal{F}_{\mathcal{W}}[t]\subset\mathcal{F}$ whose members lack sufficient connections to satisfy the robustness conditions. However, the remaining followers in $\mathcal{F}_{\mathcal{R}}[t]=\mathcal{F}\setminus\mathcal{F}_{\mathcal{W}}[t]$ (if nonempty) are indeed sufficiently connected. In other words, the subgraph induced by a set $\mathcal{L}\cup\mathcal{F}_{\mathcal{R}}[t]\subseteq\mathcal{V}$ is strongly $r$-robust with respect to $\mathcal{L}$. Using this observation, we propose a novel distributed algorithm, which we refer to as the BP-MSR algorithm.
The key idea of the BP-MSR algorithm is that agents engage in a consensus update only at time $t$ when they can locally verify that they belong to strongly $r$-robust subgraphs. For each $t\geq\mathbb{Z}_{\geq 0}$, all agents $i\in\mathcal{M}$ first compute their activation states $q_{i}[t]$ synchronously via BP (1) with $r=2F+1$ (lines 3-8). Then, leaders $i\in\mathcal{L}$ transmit their states and update them using $f_{r}$ (lines 9-11), while followers $i\in\mathcal{F}$ use $q_{i}[t]$ to locally decide whether to engage in consensus protocol synchronously (lines 12-21). If $q_{i}[t]=1$, i.e., followers $i$ belong to a strongly $(2F+1)$-robust subgraph with respect to $\mathcal{L}\cup\mathcal{A}$ at time $t$ (shown in Lemma 2), they transmit $x_{i}[t]$ and update $x_{i}[t+1]$ via the W-MSR algorithm; otherwise, they remain inactive for that time.
@@ IV-A Analysis of the BP-MSR Algorithm
By definition of $\mathcal{F}_{\mathcal{R}}[t]$, for all $i\in\mathcal{F}_{\mathcal{R}}[t]$ the following inequality must hold: $\sum_{j\in\mathcal{N}_{i}[t]\setminus\mathcal{A}}q_{j}^{i}[t,f-1]+\sum_{a\in\mathcal{A}}q_{a}^{i}[t,f-1]\geq 2F+1$. Because $q_{a}^{i}[t,k]\in\{0,1\}$ for all $a\in\mathcal{A}$, $i\in\mathcal{M}$, and $k\in[0,f]$ by 1, setting $q_{a}^{i}[t,k]=1$ for all $a\in\mathcal{A}$, $i\in\mathcal{M}$, and $k\in[0,f]$ would still ensure the inequality. Note setting $q_{a}^{i}[t,k]=1$ for all $i\in\mathcal{M}$ and $k\in[0,f]$ is equivalent to making an adversary $a\in\mathcal{A}$ a leader at time $t$. Then, applying Lemma 1 with the threshold $r=2F+1$ and initial set $\mathcal{L}\cup\mathcal{A}$, $\mathcal{G}_{\mathcal{R}}[t]$ is strongly $(2F+1)$-robust with respect to $\mathcal{L}\cup\mathcal{A}$. ∎
@@
First Claim: Here we show that $x_{i}[t]\in[m[t_{C}],M[t_{C}]]$ $\forall i\in\mathcal{F}_{\mathcal{M}}$ $\forall t\geq t_{C}$. Let $\mathcal{F}_{\mathcal{R}}[t]=\{i\in\mathcal{F}_{\mathcal{M}}\mid q_{i}[t]=1\}$. Note at each time $t\geq t_{C}$, only $i\in\mathcal{F}_{\mathcal{R}}[t]$ updates its value $x_{i}[t+1]$ with their in-neighbors’ states. If $i\notin\mathcal{F}_{\mathcal{R}}[t]$, then $x_{i}[t+1]=x_{i}[t]\in[m[t],M[t]]$. Now consider the case where $|\mathcal{F}_{\mathcal{R}}[t]|\neq 0$ and $i\in\mathcal{F}_{\mathcal{R}}[t]$. Let $\mathcal{G}_{\mathcal{R}}[t]=(\mathcal{V}_{\mathcal{R}}[t],\mathcal{E}_{\mathcal{R}}[t])$ be a subgraph of $\mathcal{G}[t]$ induced by $\mathcal{V}_{\mathcal{R}}[t]=\mathcal{A}\cup\mathcal{L}\cup\mathcal{F}_{\mathcal{R}}[t]$, which is strongly $(2F+1)$-robust with respect to $\mathcal{L}\cup\mathcal{A}$ by Lemma 2. Also, because follower $i\in\mathcal{F}_{\mathcal{R}}[t]$ has $|\mathcal{N}_{i}[t]\cap\mathcal{A}|\leq F$ by the definition of $F$-local model, $|\mathcal{N}_{i}[t]\cap\mathcal{F}_{\mathcal{R}}[t]\setminus\mathcal{A}|\geq F+1$. Then if an adversary $a\in\mathcal{A}$ has a value $x_{a}^{i}[t]>M[t]$ or $x_{a}^{i}[t]<m[t]$, it will be filtered out, since it is one of the highest or lowest values in $\mathcal{H}_{i}[t]$ or $\mathcal{O}_{i}[t]$. Hence, all values outside the interval $[m[t],M[t]]$ will be excluded from the update. Then, follower $i$ performs a convex combination of values within $[m[t],M[t]]$ to update $x_{i}[t+1]$, which implies $x_{i}[t+1]\in[m[t],M[t]]$. Lastly, $x_{i}[t]=C_{r}\in[m[t],M[t]]$ for all $i\in\mathcal{L}_{\mathcal{M}}$ and $t\geq t_{C}$. Since $x_{i}[t+1]\in[m[t],M[t]]$ for all $i\in\mathcal{M}$ and $t\geq t_{C}$, $x_{i}[t+1]\in[m[t+1],M[t+1]]\subseteq[m[t],M[t]]\subseteq\cdots\subseteq[m[t_{C}],M[t_{C}]]$.
Second Claim: [Monotonicity of Convergent Followers]
@@
Given a sequence $\mathbb{G}$, $\mathcal{F}_{\mathcal{C}}$ is fixed and known. Let $\mathcal{F}_{\mathcal{C}}^{\prime}=\mathcal{F}_{\mathcal{M}}\setminus\mathcal{F}_{\mathcal{C}}$ and $\mathcal{M}_{\mathcal{C}}=\mathcal{M}\setminus\mathcal{F}_{\mathcal{C}}^{\prime}$. Then, $\forall i\in\mathcal{F}_{\mathcal{C}}^{\prime}\ \exists\bar{t}_{i}\geq t_{C}$ such that $q_{i}[t]=0\ \forall t\geq\bar{t}_{i}$. Now we define $\overline{m}[t]=\min_{i\in\mathcal{M}_{\mathcal{C}}}x_{i}[t]$ and $\overline{M}[t]=\max_{i\in\mathcal{M}_{\mathcal{C}}}x_{i}[t]$ as the minimum and maximum states of normal agents in $\mathcal{M}_{\mathcal{C}}$ at time $t$. We first show that there exists a finite time $\bar{t}\geq t_{C}$ such that $x_{i}[t]\in[\overline{m}[\bar{t}],\overline{M}[\bar{t}]]$ for all $i\in\mathcal{F}_{\mathcal{C}}$ and $\forall t\geq\bar{t}$. Let $\bar{t}=\max_{i\in\mathcal{F}_{\mathcal{C}}^{\prime}}t_{i}<\infty$ be the time where $q_{i}[t]=0$ $\forall i\in\mathcal{F}_{\mathcal{C}}^{\prime}$ $\forall t\geq\bar{t}$. By definition, all followers $i\in\mathcal{F}_{\mathcal{C}}^{\prime}$ do not transmit their states to its out-neighbors for all $t\geq\bar{t}$, which implies removing the nodes in $\mathcal{F}_{\mathcal{C}}^{\prime}$ from the graph $\mathcal{G}[t]$ for $t\geq\bar{t}$ does not affect our analysis on the states of $\mathcal{F}_{\mathcal{C}}$. Then, applying the same reasoning as in the proof for the first claim, we have $x_{i}[t]\in[\overline{m}[t],\overline{M}[t]]\subseteq\cdots\subseteq[\overline{m}[\bar{t}],\overline{M}[\bar{t}]]$ for all $i\in\mathcal{F}_{\mathcal{C}}$ and $t\geq\bar{t}$.
@@
We now show that $|\mathcal{S}_{X}(t,\underline{\epsilon},\overline{\epsilon})|$ decreases as $t\to\infty$ with some choices of $\underline{\epsilon}$ and $\overline{\epsilon}$.

### v2

@@ I Introduction
Consensus is a process where multiple agents agree to a common value. However, standard consensus protocols are vulnerable to adversarial agents that transmit false information to the network, potentially causing significant performance degradation or even failure. This has motivated extensive research on resilient consensus, which ensures consensus in the presence of adversarial agents [8, 29, 17, 23, 10, 26, 16, 27, 9, akgün2025multi, 20, 12].
Many works rely on Mean Subsequence Reduced (MSR)-type algorithms [24, 8, 17, 27]. In the algorithms, each non-adversarial (normal) agent discards some extreme values from its neighbors before updating its state, ensuring that the updated value is not influenced by the adversarial agents. For many MSR-type algorithms to succeed, the communication graph is assumed to satisfy topological conditions called r-robustness or $(r,s)$-robustness, which are sufficient (and necessary) conditions for normal agents to achieve consensus within the convex hull of their initial values [8, 29].
A closely related problem is that of resilient leader-follower consensus, where a subset of agents (leaders) propagate a reference signal that the rest of the network (followers) aims to track, despite the presence of adversarial agents [15, 23, 2]. Resilient broadcasting problems have also been studied [6, 21, 29], where algorithms such as the Certified Propagation Algorithm (CPA) were proposed to enable a trustworthy leader to reliably broadcast information despite the presence of faulty nodes. In [7], the authors addressed the problem of resilient distributed estimation by leveraging reliable agents that are directly connected to others. Additionally, [14] established conditions under which information can be resiliently transmitted from a set of source nodes to other nodes without direct access to the information.
@@
Despite these efforts, the system behavior under MSR algorithms in non-ideal settings — where these robustness conditions of the entire network fail — remains poorly understood. Since robustness is often hard to guarantee in energy- or communication-constrained systems, analyzing these scenarios allows us to formally quantify performance degradation and limits. Some progress has been made for leaderless settings. For example,[3, 19] study community or cluster consensus, where multiple sets of agents converge to multiple distinct values when the robustness of the entire network is not sufficient. However, they also assume certain robustness within communities (clusters) and do not study cases where these conditions fail. In another line of work,[5] investigates the non-resiliency of a graph by examining the number of possible non-convergent nodes (agents that fail to reach consensus with any other nodes) when the network lacks sufficient robustness, but these results are limited to specific classes of graphs.
@@ II Preliminaries
We denote the cardinality of a set $\mathcal{S}$ as $|\mathcal{S}|$. We denote the set of non-negative and positive integers as $\mathbb{Z}_{\geq 0}$ and $\mathbb{Z}_{>0}$. We denote a simple time-varying digraph as $\mathcal{G}[t]=(\mathcal{V},\mathcal{E}[t])$ where $\mathcal{V}$ and $\mathcal{E}[t]$ are the finite vertex set and the time-varying directed edge set, respectively. We also denote $\mathbb{G}=(\mathcal{G}[t])_{t\in\mathbb{Z}_{\geq 0}}$ as a sequence of time-varying graphs. Within the set $\mathcal{V}$, we have leaders $\mathcal{L}=\{1,\dots,l\}\subset\mathcal{V}$ that propagate the reference state according to the same function $f_{r}:\mathbb{Z}_{\geq 0}\to\mathbb{R}$ to followers $\mathcal{F}=\{l+1,\dots,n\}=\mathcal{V}\setminus\mathcal{L}$. We denote $|\mathcal{L}|=l$ and $|\mathcal{F}|=f$. A directed edge $(i,j)\in\mathcal{E}[t]$ indicates that agent $j$ is able to receive information from agent $i$ at time $t$. This implies agent $i$ is an in-neighbor of agent $j$, and agent $j$ is an out-neighbor of agent $i$. The in-neighbor and out-neighbor sets of agent $i$ are denoted as $\mathcal{N}_{i}[t]=\{j\in\mathcal{V}\;|\;(j,i)\in\mathcal{E}[t]\}$ and $\mathcal{N}_{i}^{o}[t]=\{j\in\mathcal{V}\;|\;(i,j)\in\mathcal{E}[t]\}$, respectively. The extended in-neighbor set is given as $\mathcal{B}_{i}[t]=\mathcal{N}_{i}[t]\cup\{i\}$.
@@ II-A Bootstrap Percolation
At a fixed time step $t\in\mathbb{Z}_{\geq 0}$, bootstrap percolation (BP) models the spread of activation of nodes of $\mathcal{G}[t]$ from a set of initial active nodes $\mathcal{L}\subset\mathcal{V}$, given a user-defined threshold $r\in\mathbb{Z}_{>0}$ [4]. In BP, each node is either active or inactive. At each iteration of BP, each follower $i\in\mathcal{F}$ becomes active if it has at least $r$ active neighbors and remains active until the process terminates, which happens after at most $f$ iterations.
@@
$q_{i}[t,k]=\begin{cases}1&\text{if $q_{i}[t,k-1]=1$}\text{ or }\text{$\sum_{j\in\mathcal{N}_{i}[t]}q_{j}^{i}[t,k-1]\geq r$},\\ 0&\text{otherwise},\end{cases}$ (1)
@@ II-B Resilient Consensus
We review resilient leader-follower consensus when agents cannot distinguish between leaders and followers. Let $x_{i}[t]\in\mathbb{R}$ denote the consensus state of agent $i$ at time $t$, and $x_{i}^{j}[t]$ the copy received by agent $j$. Each agent $i$ shares $x_{i}[t]$ with its out-neighbors and updates its state using
@@
An agent $a\in\mathcal{V}$ is an adversary if it transmits consensus state to all out-neighbors at all $t\geq\mathbb{Z}_{\geq 0}$ but does not follow the update protocols (1) and/or (2), or if it does not send the same values to all out-neighbors at some time $t$.
The adversarial agents in this paper are a slight modification of the well-known Byzantine agents [8, 23]: each $a\in\mathcal{A}$ may manipulate both its consensus and activation states. Both followers and leaders can be adversaries. Non-adversarial agents are called normal agents. We denote $\mathcal{A}$ as the set of adversaries and $\mathcal{M}=\mathcal{V}\setminus\mathcal{A}$ as the set of normal agents, with $\mathcal{F}_{\mathcal{M}}$$\mathcal{F}\cap\mathcal{M}\neq\emptyset$ and $\mathcal{L}_{\mathcal{M}}$$\mathcal{L}\cap\mathcal{M}\neq\emptyset$.
@@
For all $a\in\mathcal{A}$, $u\in\mathcal{N}_{a}^{o}[t]$, $k\in\{0,\dots,f\}$, and $t\in\mathbb{Z}_{\geq 0}$, we have $q_{a}^{u}[t,k]\in\{0,1\}$.
Hence, an adversary cannot transmit non-binary values during the BP process. This is a direct violation of (1) and would lead to detections, which adversaries would try to avoid. In addition, we consider the following scope of threat:
@@
Lemma 1 ([11, Corollary 1 (revised)]).
@@ III Problem Statement
Although adversaries may manipulate both their consensus and activation states, our focus is on the convergence of the consensus states. That is, we do not require the activation states received by a normal agent to be correct. Nevertheless, as we demonstrate in later sections, how adversaries share their activation states influences which followers ultimately achieve convergence (see Section IV-B).
@@
Let $\mathbb{G}=(\mathcal{G}[t])_{t\in\mathbb{Z}_{\geq 0}}$ be an arbitrary sequence of time-varying digraphs under an $F$-local $\mathcal{A}$. We aim to characterize sufficient conditions under which followers in $\mathcal{F}_{\mathcal{C}}\subseteq\mathcal{F}_{\mathcal{M}}$ achieve partial resilient leader–follower consensus.
@@ IV Partial Resilient Leader-Follower Consensus
Consider a network $\mathcal{G}[t]=(\mathcal{V},\mathcal{E}[t])$ that fails to be strongly $r$-robust at time $t$. Since robustness conditions are defined over the entire network topology, that implies there exists a nonempty subset $\mathcal{F}_{\mathcal{W}}[t]\subset\mathcal{F}$ whose members lack sufficient connections to satisfy the robustness conditions. However, the remaining followers in $\mathcal{F}_{\mathcal{R}}[t]=\mathcal{F}\setminus\mathcal{F}_{\mathcal{W}}[t]$ (if nonempty) are indeed sufficiently connected. In other words, the subgraph induced by a set $\mathcal{L}\cup\mathcal{F}_{\mathcal{R}}[t]\subseteq\mathcal{V}$ is strongly $r$-robust with respect to $\mathcal{L}$. Using this observation, we propose a novel distributed algorithm, which we refer to as the BP-MSR algorithm (Algorithm 1).
The key idea of the BP-MSR algorithm is that agents engage in a consensus update only at time $t$ when they can locally verify that they belong to strongly $r$-robust subgraphs. For each $t\in\mathbb{Z}_{\geq 0}$, all agents $i\in\mathcal{M}$ first compute their activation states $q_{i}[t]$ synchronously via BP (1) with $r=2F+1$ (lines 3-8). Then, leaders $i\in\mathcal{L}$ transmit their states and update them using $f_{r}$ (lines 9-11), while followers $i\in\mathcal{F}$ use $q_{i}[t]$ to locally decide whether to engage in consensus protocol synchronously (lines 12-21). If $q_{i}[t]=1$, i.e., followers $i$ belong to a strongly $(2F+1)$-robust subgraph with respect to $\mathcal{L}\cup\mathcal{A}$ at time $t$ (shown in Lemma 2), they transmit $x_{i}[t]$ and update $x_{i}[t+1]$ via the W-MSR algorithm; otherwise, they remain inactive for that time.
@@ IV-A Analysis of the BP-MSR Algorithm
By definition, for all $i\in\mathcal{F}_{\mathcal{R}}[t]$ there exists $\bar{k}\in\{0,\dots,f-1\}$ such that the following inequality holds: $\sum_{j\in\mathcal{N}_{i}[t]\setminus\mathcal{A}}q_{j}^{i}[t,\bar{k}]+\sum_{a\in\mathcal{A}}q_{a}^{i}[t,\bar{k}]\geq 2F+1$. Because $q_{a}^{i}[t,k]\in\{0,1\}$ for all $a\in\mathcal{A}$, $i\in\mathcal{M}$, and $k\in\{0,\dots,f\}$ by 1, setting $q_{a}^{i}[t,k]=1$ for all $a\in\mathcal{A}$, $i\in\mathcal{M}$, and $k\in\{0,\dots,f-1\}$ would still ensure that the inequality holds for the same $\bar{k}$. Note setting $q_{a}^{i}[t,k]=1$ for all $i\in\mathcal{M}$ and $k\in\{0,\dots,f-1\}$ is equivalent to making an adversary $a\in\mathcal{A}$ a leader at time $t$. Then, applying Lemma 1 with the threshold $r=2F+1$ and initial set $\mathcal{L}\cup\mathcal{A}$, $\mathcal{G}_{\mathcal{R}}[t]$ is strongly $(2F+1)$-robust with respect to $\mathcal{L}\cup\mathcal{A}$. ∎
@@
First Claim: Let $\mathcal{F}_{\mathcal{R}}[t]=\{i\in\mathcal{F}_{\mathcal{M}}\mid q_{i}[t]=1\}$. Note at each time $t\geq t_{C}$, only $i\in\mathcal{F}_{\mathcal{R}}[t]$ updates its value $x_{i}[t+1]$ with their in-neighbors’ states. If $i\notin\mathcal{F}_{\mathcal{R}}[t]$, then $x_{i}[t+1]=x_{i}[t]\in[m[t],M[t]]$. Now consider the case where $|\mathcal{F}_{\mathcal{R}}[t]|\neq 0$ and $i\in\mathcal{F}_{\mathcal{R}}[t]$. Let $\mathcal{G}_{\mathcal{R}}[t]=(\mathcal{V}_{\mathcal{R}}[t],\mathcal{E}_{\mathcal{R}}[t])$ be a subgraph of $\mathcal{G}[t]$ induced by $\mathcal{V}_{\mathcal{R}}[t]=\mathcal{A}\cup\mathcal{L}\cup\mathcal{F}_{\mathcal{R}}[t]$, which is strongly $(2F+1)$-robust with respect to $\mathcal{L}\cup\mathcal{A}$ by Lemma 2. Also, because follower $i\in\mathcal{F}_{\mathcal{R}}[t]$ has $|\mathcal{N}_{i}[t]\cap\mathcal{A}|\leq F$ by the definition of $F$-local model, $|\mathcal{N}_{i}[t]\cap\mathcal{F}_{\mathcal{R}}[t]\setminus\mathcal{A}|\geq F+1$. Then if an adversary $a\in\mathcal{A}$ has a value $x_{a}^{i}[t]>M[t]$ or $x_{a}^{i}[t]<m[t]$, it will be filtered out, since it is one of the highest or lowest values in $\mathcal{H}_{i}[t]$ or $\mathcal{O}_{i}[t]$. Hence, all values outside the interval $[m[t],M[t]]$ will be excluded from the update. Then, follower $i$ performs a convex combination of values within $[m[t],M[t]]$ to update $x_{i}[t+1]$, which implies $x_{i}[t+1]\in[m[t],M[t]]$. Lastly, $x_{i}[t]=C_{r}\in[m[t],M[t]]$ for all $i\in\mathcal{L}_{\mathcal{M}}$ and $t\geq t_{C}$. Since $x_{i}[t+1]\in[m[t],M[t]]$ for all $i\in\mathcal{M}$ and $t\geq t_{C}$, $x_{i}[t+1]\in[m[t+1],M[t+1]]\subseteq[m[t],M[t]]\subseteq\cdots\subseteq[m[t_{C}],M[t_{C}]]$.
@@
Second Claim: [Monotonicity of Convergent Followers] Let $\mathcal{F}_{\mathcal{C}}^{\prime}=\mathcal{F}_{\mathcal{M}}\setminus\mathcal{F}_{\mathcal{C}}$ and $\mathcal{M}_{\mathcal{C}}=\mathcal{L}_{\mathcal{M}}\cup\mathcal{F}_{\mathcal{C}}$. Then, $\forall i\in\mathcal{F}_{\mathcal{C}}^{\prime}\ \exists\bar{t}_{i}\geq t_{C}$ such that $q_{i}[t]=0\ \forall t\geq\bar{t}_{i}$. Now we define $\overline{m}[t]=\min_{i\in\mathcal{M}_{\mathcal{C}}}x_{i}[t]$ and $\overline{M}[t]=\max_{i\in\mathcal{M}_{\mathcal{C}}}x_{i}[t]$ as the minimum and maximum states of normal agents in $\mathcal{M}_{\mathcal{C}}$ at time $t$. We first show that there exists a finite time $\bar{t}\geq t_{C}$ such that $x_{i}[t]\in[\overline{m}[\bar{t}],\overline{M}[\bar{t}]]$ for all $i\in\mathcal{F}_{\mathcal{C}}$ and $\forall t\geq\bar{t}$. Let $\bar{t}=\max_{i\in\mathcal{F}_{\mathcal{C}}^{\prime}}t_{i}<\infty$ be the time where $q_{i}[t]=0$ $\forall i\in\mathcal{F}_{\mathcal{C}}^{\prime}$ $\forall t\geq\bar{t}$. By definition, all followers $i\in\mathcal{F}_{\mathcal{C}}^{\prime}$ do not transmit their states to out-neighbors $\forall t\geq\bar{t}$. Thus, removing nodes in $\mathcal{F}_{\mathcal{C}}^{\prime}$ from the graph $\mathcal{G}[t]$ for all $t\geq\bar{t}$ does not affect our analysis on the states of $\mathcal{F}_{\mathcal{C}}$. Then, by the same reasoning as in the proof for the first claim, we have $x_{i}[t]\in[\overline{m}[t],\overline{M}[t]]\subseteq\cdots\subseteq[\overline{m}[\bar{t}],\overline{M}[\bar{t}]]$ for all $i\in\mathcal{F}_{\mathcal{C}}$ and $t\geq\bar{t}$.
@@
The set $\mathcal{S}_{X}(t,\underline{\epsilon},\overline{\epsilon})$ denotes a set of convergent followers whose states are outside of $\overline{m}[t]+\underline{\epsilon},\overline{M}[t]-\bar{\epsilon}]$. For the rest of the proof, we show that $|\mathcal{S}_{X}(t,\underline{\epsilon},\overline{\epsilon})|$ decreases as $t\to\infty$ with some suitable values of $\underline{\epsilon},\overline{\epsilon}\geq 0$. Then, we show $\overline{M}[t]-\overline{m}[t]\to 0$ as $t\to\infty$.